# Harmonic covers of skeleta

### v1

@@ 4. Metrisation of $\mathbb{Z}$-PL spaces
In this section we observe that any $\mathbb{Z}$-PL space $\Sigma$ is endowed with a canonical piecewise Lebesgue measure $\mu_{\Sigma}$, that we call the conformal measure, such that isometries are isomorphisms. We give explicit descriptions of $\mu_{\Sigma}$ in the case where $\Sigma$ is a dual complex.
@@
Let $\sigma$ be a polyhedron; recall that by convention $\sigma$ carries a $\mathbb{Z}$-affine structure: $\sigma=(|\sigma|,N_{\sigma})$ where $N_{\sigma}$ is the lattice of $\mathbb{Z}$-affine functions. For any polyhedron $\sigma$ let $H_{0}$ be the real vector space spanned by the differences of elements of $\sigma$. Then $H_{0}$ contains the maximal lattice $N_{\sigma}\cap H_{0}$ and therefore $\sigma$ inherits by translation an integral structure. For many purposes this is already a good procedure to equip $\sigma$ with a Lebesgue measure which only depends on the $\mathbb{Z}$-affine structure, it is for instance used in the works [CD12, JN20]. However for our purposes it is slightly more natural to work with a slightly different construction that can distinguish isomorphism types of polyhedra: see Example 4.6 below.
@@
Suppose $A\in\mathbb{Z}^{n\times m}$ and
@@
where $\rho\sigma=\{\rho x:x\in\sigma\}$ is the $\rho$-scaling of $\sigma$. Then $H_{\rho\sigma}\cap N$ is a full-rank lattice in $H_{\rho\sigma}$ by Lemma 4.3 (i). Let $\mu_{H_{\rho\sigma}}$ be the unique Lebesgue measure on $H_{\rho\sigma}$ such that $N\cap H_{\rho\sigma}$ has covolume $1$. We write $\mu_{\rho\sigma}$ for the restriction to $\rho\sigma$. Then we define the conformal measure on $\sigma$ as
@@
If $0\in\sigma$ we define $\mu_{\sigma}$ as the restriction of the unique Lebesgue measure on $(N_{\sigma})_{\mathbb{R}}$ such that $N_{\sigma}$ has covolume $1$.
We will call $\mu_{\sigma}$ the conformal measure on $\sigma$. The name is justified by Lemma 4.7 below.
@@
For another example, the line segment $J\subset\mathbb{R}^{n}$ given by $x+2y=1$ and $x,y\geq 0$ has length $1/2$. Note that the map $\phi:(x,y)\mapsto(2x,y)$ sends $I$ to $J$, and rescales the length by a dilation factor $2$, which is precisely the determinant of $\phi$. This will be generalised in Lemma 4.7 below and motivates our choice of the exponent $\dim\sigma+1$ in Equation (4.5). Indeed, if instead we had chosen the perhaps more natural-looking exponent $\dim\sigma$, then in the above we obtain something that we do not want: $I\to J$ would be an isometry, but $\phi$ is not unimodular (i.e. of determinant 1), and so not an isomorphism.
@@
In particular $\phi$ is a $\mathbb{Z}$-PL isomorphism (i.e. a unimodular map) if and only if it is an isometry with respect to the conformal measure, and it is strictly expanding otherwise.
@@
We can assume $0\not\in\sigma$. Let $\rho$ (resp. $\rho^{\prime}$) be the minimal $\rho\in\mathbb{Q}_{>0}$ such that $H_{\rho\sigma}\cap N_{\sigma}\neq\emptyset$ (resp. $H_{\rho^{\prime}\sigma^{\prime}}\cap N_{\sigma^{\prime}}\neq\emptyset$. Then we have
@@
Remark 4.8.
As a corollary of 4.7 we obtain the following change of variables formula: for $\phi:\sigma^{\prime}\to\sigma$ as in the lemma and every integrable $F:\sigma\to\mathbb{R}$ we have $\int_{\sigma}F\mu=[M_{\sigma^{\prime}}:M_{\sigma}]\int_{\sigma^{\prime}}(\phi^{*}F)\mu^{\prime}.$
4.9The conformal measures $\mu_{\Sigma}$.
@@
4.10.
For the remainder of this section we explicitise the construction of $\mu_{\Sigma}$ in the case where $\Sigma=D({X}^{\dagger})$ is a dual complex. As we will see the conformal measure recovers the integral Lebesgue measures of Jonsson-Nicaise [JN20, §5.1] up to a suitable normalisation by the root index, which plays the role of the scaling constant $\rho$, see Remark 4.26.
4.11Notation.
We recall (and simplify) some notation from Section 3: for $x\in F({X}^{\dagger})$, we write $\mathscr{C}_{x}\coloneqq\mathscr{C}_{F({X}^{\dagger}),x}=\mathscr{M}_{{X},x}^{\#}$ for the characteristic monoid, $M_{x}=\mathscr{C}_{x}^{\mathrm{gp}}$ for the characteristic lattice
@@
for the polyhedron of $\mathcal{D}({X}^{\dagger})$ associated to $x$. Note $\dim\sigma_{x}=r(x)-1$ and $M_{\sigma_{x}}=M_{x}$. For any $\rho\in\mathbb{R}$ we write
$H_{\rho,x}:[\varpi]\mapsto\rho$
for the affine hyperplane in $(N_{x})_{\mathbb{R}}$, so by definition $\sigma_{x}=\sigma_{\mathscr{C}_{x}}\cap H_{1,x}$. We write $N_{\rho,x}=N_{x}\cap H_{\rho,x}$. See Figures 4.17 and 4.18 for pictures. Observe the following properties:
@@
    The hyperplane $H_{\rho,x}$ is $\mathbb{Z}$-affine (with respect to $N_{x}$) if and only if $\rho\in\mathbb{Z}$.
4.12Root index.
The root index $\rho_{x}$ at $x$ is defined as the largest positive integer $\rho$ such that the image $[\varpi]$ of $\varpi$ in $\mathscr{C}_{x}$ (equivalently, in $M_{x}$) is divisible by $\rho$. It was earlier introduced in [BN20, §2].
@@
    The cone $\mathrm{rec}(\sigma_{x})$ is a rational polyhedral cone with associated lattice $N_{0,x}\cap\mathbb{R}|\mathrm{rec}(\sigma_{x})|$.
@@
4.15Rays of $C(x)$.
Let $x\in F(X^{\dagger})$. Let $r\in C(x)(1)$ be a ray of $C(x)$. Then $r$ corresponds to a unique component $E_{r}$ of $D_{{X}}$ containing $E_{x}$ by Lemma 2.9 (i). We have two cases:
@@
    Or $r$ defines a unique vertex $v_{r}\coloneqq H_{0,x}\cap r\in\sigma_{x}(0)$ and $E_{r}$ is a horizontal component (i.e. not a vertical component, equivalently $E_{r}$ is flat over $k^{\circ}$). We also write $E_{v_{r}}=E_{r}$. Note that the multiplicity $m_{v_{r}}\coloneqq\mathrm{mult}_{E_{r}}{X}_{s}$ of $E_{r}$ equals $\langle[\varpi],n_{r}\rangle$, where $n_{r}$ is the primitive ray generator of $r$, because $\mathscr{C}_{E_{r}}\cong\mathbb{N}$ is generated by the class of a local equation of $E_{r}$.
4.16.
@@
and the log-structure is the standard one induced by $\varpi$. A computation shows that $\mathrm{Spec}\,R$ is of dimension $2$. The special fiber consists of $2$ components given by $(s,t,\varpi)$ and $(u,v,\varpi)$ and these intersect in the singular point $x=(s,t,u,v,\varpi)$. The monoid $\mathscr{C}_{x}$ at $x$ is isomorphic to the monoid generated by $[s]=(1,2),[t]=(1,1),[\varpi]=(1,0),[u]=(1,-1),[v]=(1,-2)$ in $\mathbb{Z}^{2}$, in particular ${X}^{\dagger}$ is log-regular at $x$. The dual cone $C(x)$ is isomorphic to the cone in $\mathbb{R}^{2}$ generated by $[s]^{\vee}=(2,-1)$ and $[v]^{\vee}=(2,1)$. The affine line $H_{1,x}$ is given by $y=1$. It follows that $\rho_{x}=1$ and $m_{1}=(2,-1)\cdot(1,0)=2,m_{2}=(2,1)\cdot(1,0)=2$, so that $\rho_{x}=1\neq 2=\gcd(m_{1},m_{2})$.
@@
By construction, $\sigma_{x}^{-}$ is a bounded polytope and we have $\sigma_{x}^{-}\subset\sigma_{x}$ with equality if and only if $\sigma_{x}$ is bounded. Informally speaking, $\sigma_{x}^{-}$ is obtained by “cutting” $\sigma_{x}$ with the unique hyperplane parallel to $v_{1},\dots,v_{b(x)}$ which passes through $n_{b(x)+1},\dots,n_{s(x)}$
@@
coincide with the vectors $n_{i}^{\prime}\coloneqq\rho n_{i}/m_{i}\in N_{\rho,x}$ for $i=1,\dots,b(x)$. Note that $\sigma_{\rho,x}^{-}$ is obtained by scaling $\sigma_{x}^{-}$ along the cone $C(x)$ with scaling factor $\rho$. Hence we compute that
@@
Remark 4.26.
If $\lambda_{x}$ denotes the integral Lebesgue measure of Jonsson-Nicaise [JN20, §5.1] then it is immediate that $\mu_{\sigma_{x}}=\frac{1}{\rho_{x}}\lambda_{x}.$ So the conformal measure is obtained by normalising the Jonsson-Nicaise measure by the root index.
@@ 5. Specialisation of Cartier divisors
this group coincides the group of $\mathbb{Z}$-PL functions $\Sigma({X}^{\dagger})$ that are linear on the cones of $\Sigma({X}^{\dagger})$. If $\mathscr{D}$ is a Cartier divisor supported on $D_{{X}}$, then the associated $\mathbb{Z}$-PL function is denoted by $F_{\mathscr{D}}$ and is called the support function of $\mathscr{D}$. Explicitly it is described as follows: locally near a stratum $x\in F({X}^{\dagger})$ there exists a $m\in M_{x}$ such that $\mathscr{D}=\mathrm{div}\,m$, and so we have
@@
on $\sigma_{\mathscr{C}_{x}}$. Conversely we can reconstruct $\mathscr{D}$ from $[m]$ near $x$ since a choice of section $M_{x}=\mathscr{C}_{x}^{\mathrm{gp}}\to\mathscr{M}_{{X},x}^{\mathrm{gp}}$ is canonical up to a unit in $\mathscr{O}_{{X},x}^{\times}$ and so produces the same divisor $\mathrm{div}\,m$.
@@
The map $\mathscr{D}\to F_{\mathscr{D}}$ establishes a group isomorphism between Cartier divisors (resp. $\mathbb{Q}$-Cartier divisors) supported on $D_{{X}}$ and $M_{\Sigma({X}^{\dagger})}$ (resp. $(M_{\Sigma({X}^{\dagger})})_{\mathbb{Q}}$).
Moreover, effective Cartier divisors (resp. effective $\mathbb{Q}$-cartier divisors) correspond to elements of $M_{\Sigma({X}^{\dagger})}^{+}\coloneqq\lim_{x\in F({X}^{\dagger})}\mathscr{C}_{{X},x}$ (resp. $(M_{\Sigma({X}^{\dagger})}^{+})_{\mathbb{Q}_{>0}}=\lim_{x\in F({X}^{\dagger})}\mathscr{C}_{{X},x}^{\mathbb{Q}_{>0}}$).
@@
By definition, $\mathcal{D}({X}^{\dagger})$ and $\Sigma({X}^{\dagger})$ have the same sets of $\mathbb{Z}$-PL functions and so we let $M_{\mathcal{D}({X}^{\dagger})}=M_{\Sigma({X}^{\dagger})}$, this is the group of functions on $\mathcal{D}({X}^{\dagger})$ that are $\mathbb{Z}$-affine on polyhedra.
Given an element $F\in M_{\mathcal{D}({X}^{\dagger})}$ we write $\mathscr{D}_{F}$ for the associated Cartier divisor as in Lemma 5.4. Note that for all $F,F^{\prime}\in M_{\mathcal{D}({X}^{\dagger})}$ we have $\mathscr{D}_{F+F^{\prime}}=\mathscr{D}_{F}+\mathscr{D}_{F^{\prime}}$ by Lemma 5.4.
@@
Suppose $F\in M_{\mathcal{D}({X}^{\dagger})}$. Write $\mathscr{D}_{F}$ for the associated Cartier divisor as in Lemma 5.4. Then we have an equality of Weil divisors
@@
  • $E_{r}$ for the unique component of $D_{{X}}$ associated to a ray $r\in\Sigma({X}^{\dagger})(1)$ (see 4.15)
  • $n_{r}\in N_{r}\cap r$ for the primitive ray generator of a ray $r$
  • $E_{v}$ for the unique component associated to a vertex $v\in\mathcal{D}({X}^{\dagger})(0)$.
  • $m_{v}\coloneqq\mathrm{mult}_{E_{v}}{X}_{s}$ for the multiplicity of a vertex $v\in\mathcal{D}({X}^{\dagger})(0)$.
@@
Moreover, the same assertion holds for $(M_{\mathcal{D}({X}^{\dagger})})_{\mathbb{Q}}$ and $\mathbb{Q}$-Weil divisors.
@@
Let $r$ be a ray of $\Sigma({X}^{\dagger})$. Choose a local equation $\mathscr{D}_{F}=\mathrm{div}\,m$ near $E_{r}$, for some $m\in M_{r}$. Note that $M_{r}$ is a $1$-dimensional lattice dual to $n_{r}\mathbb{Z}$. By the description of 5.3 it follows $F(n_{r})=\langle m,n_{r}\rangle=\mathrm{mult}_{E_{r}}\mathscr{D}_{F}$. This gives the first equality.
For the second equality: any ray $r$ of $\Sigma({X}^{\dagger})$ is either contained in [ϖ]superscriptdelimited-[]italic-$[\varpi]^{\perp}$ or not, giving rise to a ray of the recession cone complex of $\mathcal{D}({X}^{\dagger})$ or a vertex $v\in\mathcal{D}({X}^{\dagger})$ respectively (see 4.15). For the latter case, note that $[E_{r}]=[E_{v}]$ and $F(n_{r})=F(v)\langle[\varpi],n_{r}\rangle=F(v)\mathrm{mult}_{E_{v}}\mathrm{div}\,[\varpi]=F(v)m_{v}$. ∎
@@
It is immediate to verify that the following are equivalent:

### v2

@@ 4. Metrisation of $\mathbb{Z}$-PL complexes
In this section we show that any $\mathbb{Z}$-PL complex $\Sigma$ is endowed with a canonical piecewise Lebesgue measure $\mu_{\Sigma}$, that we call the conformal measure, such that isometries are isomorphisms. We give explicit descriptions of $\mu_{\Sigma}$ in the case where $\Sigma$ is a dual complex.
@@
Let $\sigma$ be a polyhedron; recall that by convention $\sigma$ carries a $\mathbb{Z}$-affine structure: $\sigma=(|\sigma|,N_{\sigma})$ where $N_{\sigma}$ is the lattice of $\mathbb{Z}$-affine functions. Let $H_{0}$ be the real vector space spanned by the differences of elements of $\sigma$. Then $H_{0}$ contains the maximal lattice $N_{\sigma}\cap H_{0}$ and therefore $\sigma$ inherits an integral structure by translation . For many purposes this is already a good procedure to equip $\sigma$ with a Lebesgue measure which only depends on the $\mathbb{Z}$-affine structure, it is for instance used in the works [CD12, JN20]. However for our purposes it is more natural to work with a slightly different construction that can distinguish isomorphism types of polyhedra: see Example 4.6 below.
@@
Suppose $A\in\mathbb{Z}^{n\times m}$.
@@
where we write $\rho\sigma=\{\rho x:x\in\sigma\}$ for the $\rho$-scaling of $\sigma$. Then $H_{\rho\sigma}\cap N$ is a full-rank lattice in $H_{\rho\sigma}$ by Lemma 4.3 (i). Let $\mu_{H_{\rho\sigma}}$ be the unique Lebesgue measure on $H_{\rho\sigma}$ such that $N\cap H_{\rho\sigma}$ has covolume $1$. We write $\mu_{\rho\sigma}$ for the restriction to $\rho\sigma$. Then we define the conformal measure on $\sigma$ as
@@
If $0\in\sigma$ we define $\mu_{\sigma}$ as the restriction of the unique Lebesgue measure on $(N_{\sigma})_{\mathbb{R}}$ such that $N_{\sigma}$ has covolume $1$.The name “conformal measure” is justified by Lemma 4.7 below.
@@
For another example, the line segment $J\subset\mathbb{R}^{n}$ given by $x+2y=1$ and $x,y\geq 0$ has length $1/2$ with respect to $\mu_{J}$. The map $\phi:(x,y)\mapsto(2x,y)$ sends $I$ to $J$, and rescales the length by a dilation factor $2$, which is precisely the determinant of $\phi$. This will be generalised in Lemma 4.7 below and motivates our choice of the exponent $\dim\sigma+1$ in Equation (4.5). Indeed, if instead we had chosen the perhaps more natural-looking exponent $\dim\sigma$, then in the above we obtain something that we do not want: $I\to J$ would be an isometry, but $\phi$ is not unimodular (i.e. of determinant 1), so not a $\mathbb{Z}$-PL isomorphism.
@@
In particular $\phi$ is a $\mathbb{Z}$-PL isomorphism (i.e. a unimodular bijection) if and only if it is an isometry with respect to the conformal measure, and it is strictly expanding otherwise.
@@
We can assume $0\not\in\sigma$. Let $\rho$ (resp. $\rho^{\prime}$) be the minimal $\rho\in\mathbb{Q}_{>0}$ such that $H_{\rho\sigma}\cap N_{\sigma}\neq\emptyset$ (resp. $H_{\rho^{\prime}\sigma^{\prime}}\cap N_{\sigma^{\prime}}\neq\emptyset$). Then we have
@@
4.8The conformal measures $\mu_{\Sigma}$.
@@
4.9.
For the remainder of this section we explicitise the construction of $\mu_{\Sigma}$ in the case where $\Sigma=D({X}^{\dagger})$ is a dual complex of a toroidal $S$-scheme ${X}^{\dagger}$. In summary, for any $x\in F({X}^{\dagger})$ the conformal measure $\mu_{x}$ equals $\frac{1}{\rho_{x}}\lambda_{x}$ where $\lambda_{x}$ is the integral Lebesgue measure of Jonsson-Nicaise [JN20, §5.1] and the scaling constant $\rho_{x}$ is the root index (see 4.11).
4.10Notation.
Recall the notation from Section 3: for $x\in F({X}^{\dagger})$, we write $\mathscr{C}_{x}\coloneqq\mathscr{C}_{F({X}^{\dagger}),x}=\mathscr{M}_{{X},x}^{\#}$ for the characteristic monoid, $M_{x}=\mathscr{C}_{x}^{\mathrm{gp}}$ for the characteristic lattice
@@
for the polyhedron of $\mathcal{D}({X}^{\dagger})$ associated to $x$. Note $\dim\sigma_{x}=r(x)-1$ and $M_{\sigma_{x}}=M_{x}$. For any $\rho\in\mathbb{R}$ we write $H_{\rho,x}$ for the affine hyperplane in $(N_{x})_{\mathbb{R}}$ defined by $[\varpi]\mapsto\rho$, so by definition $\sigma_{x}=\sigma_{\mathscr{C}_{x}}\cap H_{1,x}$. We write $N_{\rho,x}=N_{x}\cap H_{\rho,x}$. See Figures 4.16 and 4.17 for pictures. We have the following properties:
@@
    The hyperplane $H_{\rho,x}$ is $\mathbb{Z}$-affine with respect to $N_{x}$ if and only if $\rho\in\mathbb{Z}$.
4.11Root index.
The root index $\rho_{x}$ at $x$ is defined as the largest positive integer $\rho$ such that the image $[\varpi]$ of $\varpi$ in $\mathscr{C}_{x}$ (equivalently, in $M_{x}$) is divisible by $\rho$. It was introduced in [BN20, §2].
@@
    The cone $\mathrm{rec}(\sigma_{x})$ is a rational polyhedral cone with associated lattice $N_{0,x}\cap|\mathrm{rec}(\sigma_{x})|_{\mathbb{R}}$.
@@
4.14Rays of $C(x)$.
Let $x\in F(X^{\dagger})$. Let $r\in C(x)(1)$ be a ray of $C(x)$. Then $r$ corresponds to a unique component $E_{r}$ of $D_{{X}}$ containing $E_{x}$, by Lemma 2.9 (i). We have two cases:
@@
    Or $r$ defines a unique vertex $v_{r}\coloneqq H_{0,x}\cap r\in\sigma_{x}(0)$ and $E_{r}$ is a horizontal component (i.e. not a vertical component, equivalently $E_{r}$ is flat over $k^{\circ}$). We also write $E_{v_{r}}=E_{r}$. The multiplicity $m_{v_{r}}\coloneqq\mathrm{mult}_{E_{r}}{X}_{s}$ of $E_{r}$ equals $\langle[\varpi],n_{r}\rangle$, where $n_{r}$ is the primitive ray generator of $r$, because $\mathscr{C}_{E_{r}}\cong\mathbb{N}$ is generated by the class of a local equation of $E_{r}$.
4.15.
@@
and the log-structure is the standard one induced by $\varpi$. A computation shows that $\mathrm{Spec}\,R$ is of dimension $2$ and the special fiber consists of $2$ components given by $(s,t,\varpi)$ and $(u,v,\varpi)$ and these intersect in the singular point $x=(s,t,u,v,\varpi)$. The monoid $\mathscr{C}_{x}$ at $x$ is isomorphic to the monoid generated by $[s]=(1,2),[t]=(1,1),[\varpi]=(1,0),[u]=(1,-1),[v]=(1,-2)$ in $\mathbb{Z}^{2}$, in particular ${X}^{\dagger}$ is log-regular at $x$. The dual cone $C(x)$ is isomorphic to the cone in $\mathbb{R}^{2}$ generated by $[s]^{\vee}=(2,-1)$ and $[v]^{\vee}=(2,1)$. The affine line $H_{1,x}$ is given by $y=1$. It follows that $\rho_{x}=1$ and $m_{1}=(2,-1)\cdot(1,0)=2,m_{2}=(2,1)\cdot(1,0)=2$, so that $\rho_{x}=1\neq 2=\gcd(m_{1},m_{2})$.
@@
By construction, $\sigma_{x}^{-}$ is a polytope and we have $\sigma_{x}^{-}\subset\sigma_{x}$ with equality if and only if $\sigma_{x}$ is bounded. Informally speaking $\sigma_{x}^{-}$ is obtained by “cutting” $\sigma_{x}$ with the unique hyperplane parallel to $v_{1},\dots,v_{b(x)}$ which passes through $n_{b(x)+1},\dots,n_{s(x)}$
@@
coincide with the vectors $n_{i}^{\prime}\coloneqq\rho n_{i}/m_{i}\in N_{\rho,x}$ for $i=1,\dots,b(x)$. The polytope $\sigma_{\rho,x}^{-}$ is obtained by scaling $\sigma_{x}^{-}$ along the cone $C(x)$ with scaling factor $\rho$. Hence we compute that
@@ 5. Specialisation of Cartier divisors
this group coincides with the group of $\mathbb{Z}$-PL functions $\Sigma({X}^{\dagger})$ that are linear on the cones of $\Sigma({X}^{\dagger})$. For simplicity elements of $M_{\Sigma({X}^{\dagger})}$ are also called $\mathbb{Z}$-linear functions on $\Sigma({X}^{\dagger})$. If $\mathscr{D}$ is a Cartier divisor supported on $D_{{X}}$, then the associated $\mathbb{Z}$-linearfunction is denoted by $F_{\mathscr{D}}$ and is called the support function of $\mathscr{D}$. More explicitly it is described as follows: locally near a stratum $x\in F({X}^{\dagger})$ there exists a $m\in M_{x}$ such that $\mathscr{D}=\mathrm{div}\,m$, and so we have
@@
on $\sigma_{\mathscr{C}_{x}}$. Conversely we can reconstruct $\mathscr{D}$ near $x$ from the image of $[m]$ in $M_{x}$ because a choice of section $M_{x}=\mathscr{C}_{x}^{\mathrm{gp}}\to\mathscr{M}_{{X},x}^{\mathrm{gp}}$ is canonical up to a unit in $\mathscr{O}_{{X},x}^{\times}$ and therefore produces the same divisor $\mathrm{div}\,m$.
@@
The map $\mathscr{D}\mapsto F_{\mathscr{D}}$ establishes a group isomorphism between Cartier divisors supported on $D_{{X}}$ (respectively $\mathbb{Q}$-Cartier divisors supported on $D_{{X}}$) and the group $M_{\Sigma({X}^{\dagger})}$ (respectively $(M_{\Sigma({X}^{\dagger})})_{\mathbb{Q}}$).
Moreover, effective Cartier divisors (respectively effective $\mathbb{Q}$-cartier divisors) correspond to elements of $M_{\Sigma({X}^{\dagger})}^{+}\coloneqq\lim_{x\in F({X}^{\dagger})}\mathscr{C}_{{X},x}$ (respectively $(M_{\Sigma({X}^{\dagger})}^{+})_{\mathbb{Q}_{>0}}=\lim_{x\in F({X}^{\dagger})}\mathscr{C}_{{X},x}^{\mathbb{Q}_{>0}}$).
@@
By definition $\mathcal{D}({X}^{\dagger})$ and $\Sigma({X}^{\dagger})$ have the same sets of $\mathbb{Z}$-PL functions and so we let $M_{\mathcal{D}({X}^{\dagger})}=M_{\Sigma({X}^{\dagger})}$, this is the group of functions on $\mathcal{D}({X}^{\dagger})$ that are $\mathbb{Z}$-affine on polyhedra.
Given an element $F\in M_{\mathcal{D}({X}^{\dagger})}$ we write $\mathscr{D}_{F}$ for the associated Cartier divisor as in Lemma 5.4. For all $F,F^{\prime}\in M_{\mathcal{D}({X}^{\dagger})}$ we have $\mathscr{D}_{F+F^{\prime}}=\mathscr{D}_{F}+\mathscr{D}_{F^{\prime}}$ by Lemma 5.4.
@@
Suppose $F\in M_{\mathcal{D}({X}^{\dagger})}$ and let $\mathscr{D}_{F}$ as in 5.6.Then we have an equality of Weil divisors
@@
  • $\mathrm{cyc}\mathscr{D}_{F}$ for the Weil divisor associated to $\mathscr{D}_{F}$,
  • $E_{r}$ for the unique component of $D_{{X}}$ associated to a ray $r\in\Sigma({X}^{\dagger})(1)$ (see 4.14)
  • $n_{r}\in N_{r}\cap r$ for the primitive ray generator of a ray $r\in\Sigma({X}^{\dagger})(1)$,
  • $E_{v}$ for the unique component associated to a vertex $v\in\mathcal{D}({X}^{\dagger})(0)$,
  • $m_{v}\coloneqq\mathrm{mult}_{E_{v}}{X}_{s}$ for the multiplicity of a vertex $v\in\mathcal{D}({X}^{\dagger})(0)$,
@@
In addition the same assertion holds for $(M_{\mathcal{D}({X}^{\dagger})})_{\mathbb{Q}}$ and $\mathbb{Q}$-Weil divisors.
@@
Let $r$ be a ray of $\Sigma({X}^{\dagger})$. Choose a local equation $\mathscr{D}_{F}=\mathrm{div}\,m$ near $E_{r}$, for some $m\in M_{r}$. Since $M_{r}$ is a $1$-dimensional lattice dual to $n_{r}\mathbb{Z}$, the description of 5.3 implies $F(n_{r})=\langle m,n_{r}\rangle=\mathrm{mult}_{E_{r}}\mathscr{D}_{F}$. This gives the first equality.
For the second equality: by 4.14 any ray $r$ of $\Sigma({X}^{\dagger})$ is either contained in [ϖ]superscriptdelimited-[]italic-$[\varpi]^{\perp}$ or not, giving rise to a ray of the recession cone complex of $\mathcal{D}({X}^{\dagger})$ or a vertex $v\in\mathcal{D}({X}^{\dagger})$ respectively. In the second case $[E_{r}]=[E_{v}]$ and $F(n_{r})=F(v)\langle[\varpi],n_{r}\rangle=F(v)\mathrm{mult}_{E_{v}}\mathrm{div}\,[\varpi]=F(v)m_{v}$. ∎
@@
By Lemma 5.4 the following assertions are equivalent: